What is the answere to the following question:
The line passes through (8,-3) and is parallel to the line whose equation is y=2x+3

Answer by Edwin McCravy(20055)

It has the same slope, so it's

y=2x+b

We find b by substituting (x,y) = (8,-3)

 -3=2(8)+b
 -3=16+b
-19=b

Answer: y=2x-19

Edwin